OpenStudy (anonymous):

The maximum distance the earth is from the sun is 152,000,000 Km and the minimum distance is 147,000,000 Km. Write an absolute value equation where d represents the maximum and minimum distance between earth and the sun in millions of kilometers.

OpenStudy (anonymous):

i can give you a hint... \[\huge \min \le Earth_{sun} \le \max\]

OpenStudy (anonymous):

@agentc0re |152,000,000 - 147,000,000| = 5,000,000 ?

OpenStudy (anonymous):

I just realized it wanted something in absolute values... sorry i must have read too fast.. let me re think

OpenStudy (anonymous):

\[\huge \left| x \right| \le 1\]is equal to saying \[\huge -1 \le x \le 1\] For the time being, I'm going to drop all these zeros okay? so the difference between the two is 5.

OpenStudy (anonymous):

Wow i am getting impatient I think I might just take a 0 for this question

OpenStudy (anonymous):

okay